Please help...these word problems are very humbling for me...
Average movie prices in general are lower in the United States than in other countries. It would cost $81.35 to buy three tickets in Japan plus two tickets in Switzerland. Three tickets in Switzerland plus two tickets in Japan would cost $77.05. How much does an average movie cost in each of these countries?
I really appreciate your help...thank you...

Answer by Edwin McCravy(20086)

Please help...these word problems are very humbling for me...
Average movie prices in general are lower in the United States than in other countries. It would cost $81.35 to buy three tickets in Japan plus two tickets in Switzerland. Three tickets in Switzerland plus two tickets in Japan would cost $77.05. How much does an average movie cost in each of these countries?
I really appreciate your help...thank you...

Word problems are only a matter of translating English sentences into 
Algebra sentences.

>>...It would cost $81.35 to buy three tickets in Japan plus two tickets in Switzerland...<<
That English sentence translates into this Algebra sentence:

                   3J + 2S = 81.35

>>...Three tickets in Switzerland plus two tickets in Japan would cost $77.05....<<
That English sentence translates into this Algebra sentence:

                   3S + 2J = 77.05

Solve the system of equations:

3J + 2S = 81.35
3S + 2J = 77.05

Reverse the terms on the left of one of the equations so
like letters will line up vertically, so we can solve the
system by elimination rather than substitution.

3J + 2S = 81.35
2J + 3S = 77.05

This is more easily solved by elimination since there is
no term J, -J, S or -S.  If there had been, substitution
would have been easier.

Multiply the first equation through by -2 and the second 
one through by +3

-6J - 4S = -162.70
 6J + 9S =  231.15
------------------
      5S = 68.45
       S = 13.69

Start over to find J

3J + 2S = 81.35
2J + 3S = 77.05

Multiply the first equation through by -3 and the second 
one through by 2

-9J - 6S = -244.05
 4J + 6S =  154.10
------------------
-5J      =  -89.95
       J =   17.99 

Answer: An average movie ticket costs $17.99 in Japan and $13.69 
in Switzerland.

Edwin